OpenStudy (anonymous):

The town is organizing a Fourth of July parade. There will be two sizes of floats in the parade( 30 ft and 15 ft). A space of 10 ft will be left after each float. A. The parade must be at least 150 ft long, but less than 200 ft long. What combinations of large and small floats are possible? B. Large floats cost $600 to operate. The town has a budget of $2500 to operate the floats. How does this change your answer to part A? What combinations of large and small floats are possible?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@samvech0011

jimthompson5910 (jim_thompson5910):

"There will be two sizes of floats in the parade( 30 ft and 15 ft). A space of 10 ft will be left after each float." so the 30 ft float needs 30+10 = 40 ft the 15 ft float needs 15+10 = 25 ft

jimthompson5910 (jim_thompson5910):

x = # of 30 ft floats y = # of 15 ft floats

OpenStudy (anonymous):

i need help with part b

jimthompson5910 (jim_thompson5910):

the first equation is 40x + 25y = 150 the 40x is the total amount for the x floats of 30 ft the 25y is the total amount for the y floats of 15 ft

jimthompson5910 (jim_thompson5910):

what combos did you get?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

i got it done

jimthompson5910 (jim_thompson5910):

alright